# Investigating the Fermi-Hubbard model by the Tensor-Backflow method

### v3

@@ Abstract
Recently, a variational wave-function based on the tensor representation of backflow corrections has achieved state-of-the-art energy precision in solving Fermi-Hubbard-type models tensorbackflow . However, the Fermi-Hubbard model is very challenging to solve, and the validity of a method relies on investigating the ground state’s physical property. For simplicity, we name the tensor representation of backflow corrections as the Tensor-Backflow in this work. We apply the Tensor-Backflow method to investigate the Fermi-Hubbard model on two-dimensional lattices as large as 256 sites, under various interaction strengths $U$, electron fillings $n$, next-nearest-neighbor hopping strengths $t^{\prime}$ and boundary conditions. Instead of considering backflow terms from all sites, competitive results are achieved by considering nearest-neighbor or next-nearest-neighbor backflow terms. Meanwhile the variational wave-function is not enforced on geometric symmetries. When $t^{\prime}$=0, by considering nearest-neighbor backflow terms, linear stripe order is sucessfully obtained for the case of nn=0.875 and UU=8 on the $16\times 16$ lattice under periodoc boundary condition. For a similar case under open boundary condition, obtained energy is only $4.5\times 10^{-4}$ higher than the state-of-the-art method fPEPS with the bond dimension DD=20. Comparing to state-of-the-art neural network results, energies are competitive and relative errors are below $5\times 10^{-3}$. For cases of nn=0.8 and 0.9375, results consistent with the phase diagram from AFQMC are obtained by direct optimizations. When $t^{\prime}$=-0.2, by considering next-nearest-neighbor backflow terms, obtained energies are competitive or even lower than state-of-the-art neural network results. For example, obtained energy for nn=0.875, UU=8 on the $12\times 12$ lattice under PBC is $8.1\times 10^{-4}$ lower comparing to that from the neural network state. Therefore, the Tensor-Backflow method has strong representation abilities for the Fermi-Hubbard model.
@@ I Introduction
The Fermi-Hubbard model is a simple quantum lattice model for describing correlated electrons. Solving the Fermi-Hubbard model is important for understanding the mechanism of superconductivity Hubbard_SC1 ; Hubbard_SC2 ; Hubbard_SC3 . However, investigating the Fermi-Hubbard model on a finite-sized lattice is very challenging. Because states with different orders near the ground state have energies so similar that the optimization is easily to get trapped on a local minima. For finite systems, the lattice size should be sufficiently large to reduce the boundary effect. Therefore, to identify the property of the ground state, numerical methods are required to be unbiased and scalable. To achieve this goal, many numerical methods have been developed and benchmarked Hubbard_benchmark ; Manybody_benchmark ; Hubbard_DMET ; Hubbard_DMRG ; Hubbard_NN1 ; Hubbard_NN2 ; Hubbard_NN3 ; Hubbard_NN4 ; Hubbard_TNS ; tJ_fPEPS ; Hubbard_fPEPS .
@@
For example, the density-matrix-renormalization-group (DMRG) is suitable for one-dimensional or quasi one-dimensional systems Hubbard_benchmark ; Hubbard_DMRG ; Hubbard_DMRG2 ; Hubbard_DMRG3 ; Hubbard_DMRG4 , but the accuracy is not satisfactory for two-dimensional systems. The Quantum-Monte-Carlo (QMC) has very high accuracy, but computational complexity is too high for systems within the “sign problem” QMC_sign_1 ; QMC_sign_2 ; QMC_sign_3 . The projected-entangled-pair-state (PEPS) can achieve high accuracy for two-dimensional lattices after variational optimizations tJ_fPEPS ; Hubbard_fPEPS ; PEPS_liu , however the computational complexity is high especially for periodic boundary condition (PBC). For the infinite-PEPS (iPEPS) iPEPS and the density-matrix-embedding-theory (DMET) Hubbard_DMET , the size of the supercell introduces a prior condition on the ground state.
@@
The variational-monte-carlo (VMC) has no “sign problem”, however its effectiveness depends on the variational wave-function. Recently, the neural-network based variational wave-functions have been intensively benchmarked RBM_science ; CNN_J1J2_1 ; CNN_J1J2_2 ; CNN_J1J2_3 ; J1J2_SOTA1 ; J1J2_SOTA2 ; J1J2_SOTA3 ; Hubbard_NN1 ; Hubbard_NN2 . By the deep structure, neural-networks can capture long-range correlations or entanglements in the quantum state, thus achieving state representation ability beyond the area-law entanglement NN_represent1 ; NN_represent2 . Although backflow corrections by neural-networks have obtained high accuracy in continuous systems NN_continuous1 ; NN_continuous2 ; NN_continuous3 For the lattice Fermi-Hubbard model, high optimization difficulties of neural networks lead to unsatisfied energy precision on large lattices Hubbard_NN1 ; Hubbard_NN2 or requirements of pre-conditions on wave-functions Hubbard_NN3 ; Hubbard_NN4 . Recently, the Tensor-Backflow method has achieved state-of-the-art energy precision at reasonable optimization costs for Fermi-Hubbard-type models tensorbackflow . The wave-function is based on backflow corrections backflow1 ; backflow2 , meanwhile the representation ability is significantly enhanced by the tensor representation of the wave-function.
@@
In this work, we use the Tensor-Backflow method well-developed in the previous literature tensorbackflow to investigate the Fermi-Hubbard model under various electron fillings $n$, interaction strengths $U$, next-nearest-neighbor hopping strengths $t^{\prime}$ and boundary conditions on lattices as large as 256 sites. Without prior enforcements on geometric symmetries, achieved energies are competitive or even better than current state-of-the-art results. When $t^{\prime}$=0, the ground state with a linear stripe order is achieved without applying pinning-fields for the case of nn=0.875 and UU=8 on the $16\times 16$ lattice under PBC. Periods for spin-density-wave (SDW) and charge-density-wave (CDW) are 16 and 8, respectively. For cases of nn=0.875, UU=10 and 12, linear stripe orders are preserved when transferring from the wave-function converged for UU=8, meanwhile energies are lower comparing to other orders achieved by direct optimzations. For cases of n=n=0.8 with UU=4, 8 and nn=0.9375 with UU=4, results are consistent with the phase diagram of AFQMC AFQMC_phase_diagram ; AFQMC_n0.9375 . For more challenging cases of $t^{\prime}$=-0.2 where non-zero superconductivity is expected Hubbard_SC1 , energy achieved by Tensor-Backflow is competitive or even lower than those from the state-of-the-art neural network state (NQS) Hubbard_NN3 , demonstrating the strong representation ability of Tensor-Backflow.
@@
This paper is organized as follows: Sec.II introduces the Tensor-Backflow method as well as the VMC optimization and a Lanczos optimization. Computational complexity of the method is introduced. Subsec.III.1 compares two optimization strategies. Subsec.III.2 presents energy comparisons respect to current state-of-the-art methods such as fPEPS, AFQMC, NQS, etc. Subsec.III.3 presents results for nn=0.875 with $t^{\prime}$=0 and $t^{\prime}$=-0.2, including energy comparisons, ground state patterns, charge and spin structure factors. Subsec.III.4 presents results for nn=0.8 and 0.9375, with energy results as well as charge and spin structure factors. Sec.IV concludes this paper.
@@
II.1 Wave-function definitions
In this subsection, we briefly introduce the Tensor-Backflow method developed in the previous literature tensorbackflow . An unrestricted form of backflow corrections is proposed in literature tensorbackflow : the summation for backflow terms are not only on positions but also on spins. The fictitious orbital is constructed as
$\phi^{B}_{k\sigma_{k}}(\mathbf{r}_{i,\sigma_{i}})=\phi_{k\sigma_{k}}(\mathbf{r}_{i,\sigma_{i}})+\sum_{j}c_{ij}[\mathbf{S}]\sum_{\sigma_{j}=\pm 1}\phi_{k\sigma_{k}}(\mathbf{r}_{j,\sigma_{j}}),$ (1)
where $c_{ij}[\mathbf{S}]$ is the coefficient depending on the configuration $|\mathbf{S}\rangle$ and $\phi_{k\sigma_{k}}$ is the Hatree-Fock (HF) orbital.
Based on the backflow formula, the NQS can improve the representation ability by learning backflow terms Hubbard_NN1 ; Hubbard_NN3 ; NN_continuous1 ; NN_continuous2 ; NN_continuous3 . Comparing to NQS Hubbard_NN2 ; Hubbard_NN3 , we use tensor representation of the wave-function to improve the representation ability, meanwhile keeping the optimization feasible for first-order-gradient methods tensorbackflow .
For a function with discrete variables of $x_{1}$ and $x_{2}$: $y=f(x_{1},x_{2})$, it can be represented by a tenesor $\xi$: $y=\xi_{x_{1},x_{2}}$, where $x_{1}$ and $x_{2}$ are indices in the $\xi$. We consider coefficients $c_{ij}$ in Eq.(1) depend on local configurations $\mathbf{s}(\mathbf{r}_{i})$ and $\mathbf{s}(\mathbf{r}_{j})$ instead of $|\mathbf{S}\rangle$ for simplicity backflow1 ; backflow2 . Then, the independent variables in Eq.(1) are particle position $\textbf{r}_{i}$, orbital number $k$ for $\sigma_{k}=\pm 1$, backflow indices $j$, configurations $\mathbf{s}(\mathbf{r}_{i})$ and $\mathbf{s}(\mathbf{r}_{j})$. Based on independent degrees of freedom in the wave-function, a tensor $g$ is constructed with each dimension in $g$ corresponding to the independent variable in Eq.(1) respectively. Therefore, the total dimension of $g$ is tensorbackflow :
$[M,N,d,Q,d],$ (2)
@@ II.1 Wave-function definitions
where $M$ is the site number, $N$ is the total particle number, $d$ is the degrees of freedom each site and $Q$ is the number of backflow terms including the particle itself. First and second $d$ are for $\mathbf{s}(\mathbf{r}_{i})$ and $\mathbf{s}(\mathbf{r}_{j})$, respectively. When considering backflow terms from nearest-neighbors and all sites, there is: $Q=5$ and $N$, respectively.
The wave-function is a Slater-determinant of fictitious particles:
@@
$M_{ik}^{B}=\sum_{q}g[\mathbf{r}_{i},k,\mathbf{s}(\mathbf{r}_{i}),q,\mathbf{s}(\mathbf{r}_{q})],$ (4)
where $q$ is the summation index including the backflow terms and the $i$-th particle position.
Therefore, the wave-function $w(\mathbf{S})$ is generated by a summation of elements from tensor $g$ based on the backflow rule. In the wave-function, there are no enforcements on translational or rotational symmetries, and the total number of variational parameters is the multiplication of all indices of $g$. Comparing to representing correlations by decomposed tensors CP_tensor , in this work, the representation ability is achieved by the tensor $g$ natually containing high order correlations tensorbackflow . For example, $g$ can represent high order correlations between $\mathbf{s}(\mathbf{r}_{i})$ and $\mathbf{s}(\mathbf{r}_{j})$. Such flexible representation ability of $g$ as well as the simple construction of the wave-function lead to low optimization difficulty, therefore competitive precision is achieved within the first-order-gradient method.

### v4

@@ Abstract
Recently, a variational wave-function based on the tensor representation of backflow corrections along with a Lanczos step has achieved state-of-the-art energy precision for Fermi-Hubbard-type models. However, solving the Fermi-Hubbard model remains challenging, and the validity of any method must be assessed through the ground state’s physical properties. For simplicity, we refer this method as Tensor-Backflow. In this work, we apply the Tensor-Backflow method to investigate the Fermi-Hubbard model on two-dimensional lattices up to 256 sites, exploring various interaction strengths $U$, electron fillings $n$, next-nearest-neighbor hopping $t^{\prime}$, and boundary conditions. By considering backflow terms from nearest- or next-nearest-neighbor sites, we achieve competitive results without enforcing geometric symmetries on the variational wave-function. The optimizations were stable from a prior unrestrictied Hartree-Fock state, followed by adding backflow corrections. Meanwhile, changing interaction strengths in the prior unrestrictied Hartree-Fock state is helpful to bypass the local minima. When $t^{\prime}$=0, by considering nearest-neighbor backflow terms, linear stripe order emerges successfully for the case of nn=0.875 and UU=8 on a $16\times 16$ lattice with periodic boundary conditions. In a similar case with open boundary conditions, the energy obtained is only $4.5\times 10^{-4}$ higher than the state-of-the-art method fPEPS with bond dimension DD=20. Compared to state-of-the-art neural network methods, the energies obtained using the Tensor-Backflow approach are competitive, with relative errors below $5\times 10^{-3}$. For nn=0.8 and nn=0.9375, direct optimizations yield results consistent with the phase diagram from AFQMC. When $t^{\prime}$=-0.2, considering next-nearest-neighbor backflow terms leads to energies that are either competitive with or even lower than those from state-of-the-art neural network approaches. For instance, for nn=0.875 and UU=8 on a $12\times 12$ lattice with periodic boundary conditions, the energy obtained is $8.1\times 10^{-4}$ lower than that from the neural network result. Thus, the Tensor-Backflow method demonstrates strong representational capabilities for solving the Fermi-Hubbard model.
@@ I Introduction
The Fermi-Hubbard model is a fundamental quantum lattice model used to describe correlated electrons and plays a key role in understanding the mechanisms of superconductivity Hubbard_SC1 ; Hubbard_SC2 ; Hubbard_SC3 . However, solving the Fermi-Hubbard model, particularly on finite-sized lattices, presents significant challenges. Near the ground state, the energies of states with different ordering can be extremely close, making the optimization process easy to getting trapped in local minima. To mitigate boundary effects, the lattice size must be sufficiently large when studying finite systems. As such, identifying the ground state properties requires numerical methods that are both unbiased and scalable. To this end, various numerical approaches have been developed and benchmarked for solving the Fermi-Hubbard model Hubbard_benchmark ; Manybody_benchmark ; Hubbard_DMET ; Hubbard_DMRG ; Hubbard_NN1 ; Hubbard_NN2 ; Hubbard_NN3 ; Hubbard_NN4 ; Hubbard_TNS ; tJ_fPEPS ; Hubbard_fPEPS .
@@
For instance, the density-matrix renormalization group (DMRG) is well-suited for one-dimensional or quasi-one-dimensional systems Hubbard_benchmark ; Hubbard_DMRG ; Hubbard_DMRG2 ; Hubbard_DMRG3 ; Hubbard_DMRG4 , but its accuracy is not satisfactory for two-dimensional systems. The Quantum Monte Carlo (QMC) method offers high accuracy but suffers from high computational complexity, particularly due to the “sign problem” in certain systems QMC_sign_1 ; QMC_sign_2 ; QMC_sign_3 . The projected-entangled-pair state (PEPS) can achieve high accuracy for two-dimensional lattices after variational optimization tJ_fPEPS ; Hubbard_fPEPS ; PEPS_liu , though its computational cost remains high, especially for periodic boundary conditions (PBC). Similarly, both infinite-PEPS (iPEPS) iPEPS and density-matrix embedding theory (DMET) Hubbard_DMET impose constraints on the ground state due to the size of the supercell used.
@@
The variational Monte Carlo (VMC) method avoids the “sign problem”, but its effectiveness depends critically on the choice of variational wave-function. Recently, neural-network-based variational wave-functions have been extensively benchmarked RBM_science ; CNN_J1J2_1 ; CNN_J1J2_2 ; CNN_J1J2_3 ; J1J2_SOTA1 ; J1J2_SOTA2 ; J1J2_SOTA3 ; Hubbard_NN1 ; Hubbard_NN2 . Due to their deep structure, neural networks can capture long-range correlations and entanglements in quantum states, enabling state representations that go beyond the area-law entanglement NN_represent1 ; NN_represent2 . While backflow corrections implemented via neural networks have demonstrated high accuracy in continuous systems NN_continuous1 ; NN_continuous2 ; NN_continuous3 , their application to the lattice Fermi-Hubbard model has challenges. The high optimization difficulties of neural networks lead to unsatisfactory energy precision for large lattices Hubbard_NN1 ; Hubbard_NN2 or require specific preconditions on the wave-function Hubbard_NN3 ; Hubbard_NN4 . Recently, the Tensor-Backflow along with a Lanczos step has achieved state-of-the-art for Fermi-Hubbard-type models tensorbackflow . This approach combines backflow corrections backflow1 ; backflow2 with a tensor representation of the wave-function. The optimization of the Tensor-Backflow wave-function is stable from optimizing a prior unrestrictied Hartree-Fock (UHF), followed by including backflow corrections in the wave-function.
@@
In this work, we apply the well-established Tensor-Backflow method tensorbackflow to investigate the Fermi-Hubbard model across a range of electron fillings $n$, interaction strengths $U$, next-nearest-neighbor hopping strengths $t^{\prime}$, and boundary conditions on lattices as large as 256 sites. Without enforcing geometric symmetries, the achieved energies are competitive with, or even exceed, those of current state-of-the-art methods. For $t^{\prime}$=0, we obtain a ground state with linear stripe order for nn=0.875 and UU=8 on a $16\times 16$ lattice under periodic boundary conditions (PBC), without applying any pinning fields. The periods of spin-density-wave (SDW) and charge-density-wave (CDW) are 16 and 8, respectively. For higher interaction strengths, UU=10 and UU=12, the linear stripe order is preserved when starting from the wave-function converged for UU=8, and the energies are lower than those for other orders obtained through direct optimization. For nn=0.8 with UU=4, 8 and nn=0.9375 with UU=4, the results are consistent with the phase diagram from AFQMC AFQMC_phase_diagram ; AFQMC_n0.9375 . In more challenging cases with $t^{\prime}$=-0.2, where non-zero superconductivity is expected Hubbard_SC1 , the energies achieved by Tensor-Backflow is competitive with, or even lower than, those from state-of-the-art neural network states (NQS) Hubbard_NN3 , demonstrating the strong representational power of the Tensor-Backflow approach.
@@
This paper is organized as follows: Sec. (II) introduces the Tensor-Backflow method, including the variational Monte Carlo (VMC) optimization and Lanczos step, and discusses the computational complexity of the method. Subsec. (III.1) compares two optimization strategies. Subsec. (III.2) presents energy comparisons with current state-of-the-art methods, including fPEPS, AFQMC, NQS, and others. Subsec. (III.3) presents results for nn=0.875 with both $t^{\prime}$=0 and $t^{\prime}$=-0.2, including energy comparisons, ground state patterns, and charge and spin structure factors. Subsec. (III.4) discusses results for nn=0.8 and nn=0.9375, providing energy results as well as charge and spin structure factors. Sec. (IV) concludes the paper.
@@
II.1 Variational wave-function definitions
In this subsection, we introduce the Tensor-Backflow method developed in the previous literature tensorbackflow . The backflow correction is originally derived from the transformation of the position of particles backflow1 :
$\mathbf{r}_{\alpha}^{B}=\mathbf{r}_{\alpha}+\sum_{\beta}\eta_{\alpha\beta}[\mathbf{S}](\mathbf{r}_{\beta}-\mathbf{r}_{\alpha}),$ (1)
where $\mathbf{r}_{\alpha}$ are actual particle positions and $\eta_{\alpha\beta}[\mathbf{S}]$ are variational parameters depending on the many-body state $|\mathbf{S}\rangle$. The original backflow correction is constructed by a linear combination of orbitals backflow1 ; backflow2 :
$\phi^{B}_{k\sigma_{k}}(\mathbf{r}_{i,\sigma_{i}})=\phi_{k\sigma_{k}}(\mathbf{r}_{i,\sigma_{i}})+\sum_{j}c_{ij}[\mathbf{S}]\phi_{k\sigma_{k}}(\mathbf{r}_{j,\sigma_{j}}),$ (2)
where $c_{ij}[\mathbf{S}]$ is the coefficient depending on the configuration $|\mathbf{S}\rangle$ and $\phi_{k\sigma_{k}}$ is the Hatree-Fock (HF) orbital. In the original backflow backflow1 ; backflow2 , the coefficients $c_{ij}$ in Eq.(2) are assumed to depend on local configurations $\mathbf{s}(\mathbf{r}_{i})$ and $\mathbf{s}(\mathbf{r}_{j})$, rather than on the full state $|\mathbf{S}\rangle$ for simplicity. The final wave-function is the Slater determinant: $w(\mathbf{S})=\mathrm{det}M_{ik}^{B}$, where $M_{ik}^{B}=\langle\mathbf{r}_{i}|\phi^{B}_{k\sigma_{k}}\rangle$.
The backflow corrections defined by Eq.(1) and (2) are applied only to positions. As a result, the matrix $M_{ik}^{B}$ in the Slater determinant is quasi-diagonal with respect to different spins, which limits the representation ability for Hamiltonians with spin couplings. To enhance the representation ability for spin-coupled systems, we consider an unrestricted form of backflow corrections. This is achieved by introducing backflow terms that depend not only on positions but also on spins. The orbital is constructed as tensorbackflow :
$\phi^{B}_{k\sigma_{k}}(\mathbf{r}_{i,\sigma_{i}})=\phi_{k\sigma_{k}}(\mathbf{r}_{i,\sigma_{i}})+\sum_{j}c_{ij}\sum_{\sigma_{j}=\pm 1}\phi_{k\sigma_{k}}(\mathbf{r}_{j,\sigma_{j}}).$ (3)
Here in Eq.(3), we adopt the similar assumption for the coefficient $c_{ij}$ as in the original backflow in Eq.(2).
For the state representation, the backflow corrected orbital can involve complex relationships among the orbital index $k\sigma_{k}$ and the position $\mathbf{r}_{j}$ in the HF orbital and the configurations of $\mathbf{s}(\mathbf{r}_{i})$ and $\mathbf{s}(\mathbf{r}_{j})$ in $c_{ij}$. This complexity goes beyond the representation ability of the product of $c_{ij}$ and $\phi_{k\sigma_{k}}$ Hubbard_NN1 ; Hubbard_NN3 . To achieve the enhanced representation ability based on the variables in Eq.(3), we develop a mathematical form capable of representing arbitrary functions of these variables. A straightforward approach is to represent Eq.(3) by a single high-dimensional tensor: $g_{\mathbf{r}_{i},k\sigma_{k},j,\mathbf{s}_{i},\mathbf{s}_{j}}$, where each index in the tensor corresponds to an independent variable in Eq.(3tensorbackflow . This construction ensures high representation ability for backflow corrections on lattices.
Here we construct the tensor $g$, as well as the final wave-function, by considering the independent variables in Eq.(3). The variables include the particle positions $\mathbf{r}_{i}$, the orbital index $k\sigma_{k}$ with $\sigma_{k}=\pm 1$, the backflow indices $j$, and the configurations $\mathbf{s}(\mathbf{r}_{i})$ and $\mathbf{s}(\mathbf{r}_{j})$. Based on the definition, the final dimension of $g$ is tensorbackflow :
$[M,N,Q,d,d],$ (4)
@@ II.1 Variational wave-function definitions
where $M$ denotes the number of sites, $N$ is the total number of particles, and $d$ represents the number of degrees of freedom at each site. The first and second occurrences of $d$ correspond to $\mathbf{s}(\mathbf{r}_{i})$ and $\mathbf{s}(\mathbf{r}_{j})$, respectively. In this work, we consider backflow contributions from both the site $\mathbf{r}_{i}$ and its surrounding sites. Consequently, QQ=5 when backflow terms from nearest-neighbor sites are included, and QQ=$M$ when backflow terms from all sites are considered.
The final wave-function is a Slater determinant of fictitious particles:
@@
$M_{ik}^{B}=\sum_{j}g_{\mathbf{r}_{i},k\sigma_{k},j,\mathbf{s}(\mathbf{r}_{i}),\mathbf{s}(\mathbf{r}_{j})},$ (6)
where $j$ is the summation index including the backflow terms.
Because $g$ is a single tensor, correlations between $\mathbf{s}(\mathbf{r}_{i})$ and $\mathbf{s}(\mathbf{r}_{j})$ are achieved naturally by the backflow rule. In contrast, in the literature CP_tensor , such correlations are introduced through products of decomposed tensors. Compared with tensor decompositions, the cost for using a single tensor is the large parameter number. The total number of variational parameters equals the product of the dimensions of all indices of $g$. However, only a subset of these parameters is involved in generating a $w(\mathbf{S})$.
By constructing $g$, each site is associated with a high-dimensional tensor whose indices include $k\sigma_{k}$ and $\mathbf{s}(\mathbf{r}_{i})$, as well as those of its surroundings sites $j$ and $\mathbf{s}(\mathbf{r}_{j})$. Although the procedure for generating $w(\mathbf{S})$ is very different, the similar structure of $g$ to that of a two-dimensional tensor network results in high efficiency for first-order optimization tJ_fPEPS ; PEPS_liu . In the construction of the wave-function, no constraints are imposed on translational or rotational symmetries. The flexibility of the representation ability of $g$, combined with the straightforward construction of the wave-function, leads to the effectiveness of first-order gradient methods.
To stabilize the initial optimization and reduce the computational cost, we start by optimizing a prior UHF state. The UHF is represented by a tensor: $g_{\mathbf{r}_{i},k\sigma_{k},\sigma_{i}}^{\mathrm{UHF}}$, where $\sigma_{i}$ denotes the spin of the particle at site $\mathbf{r}_{i}$. Based on the UHF, we extend the tensor by introducing additional indices to incorporate backflow corrections. The resulting tensor takes the form: $g_{\mathbf{r}_{i},k\sigma_{k},\sigma_{i},2,j,\mathbf{s}(\mathbf{r}_{j})}$ tensorbackflow , where the dimension of 2 distinguishes the double occupation at site $\mathbf{r}_{i}$. Although the energy convergence of an UHF is very stable, the use of a prior UHF may introduce bias into the variational wave function UHF . In this work, we avoid such bias when generating the results in Sec. (III) and compare the prior UHF with the converged wave-function in App.(B).